# MiniLLM: Knowledge Distillation of Large Language Models

### v4

@@ A.2 Derivation of Equation 2
We compute the gradient of $\mathcal{L}(\theta)=\operatorname{KL}[q_{\theta}||p]$ with respect to $\theta$ using the Policy Gradient Theorem [60]:
@@ B.1 Training Details
Our baselines include SFT w/o KD, KD, and SeqKD. For models with less than 1.3B parameters, we search for the learning rates in [5e-4, 1e-4, 5e-5], the batch sizes in [32, 64], and train these models for 20 epochs. For other models, we search for the learning rate in [5e-5, 1e-5, 5e-6], the batch sizes in [32, 64], and train these models for 10 epochs. For KD, we follow [63] to mix the distillation loss with the language modeling loss on the ground truth responses by a mixture rate of 0.5. The checkpoints of each baseline are selected by the Rouge-L [39] scores on the validation set because, as stated in previous works [76, 51], we also find that Rouge-L is better correlated with human judgments.
@@
As stated in Section 2.3, training of MiniLLM has two phases which is similar to Reinforcement Learning from Human Feedback (RLHF;51).
@@
  • Phase 2: We continuously train the model from Phase 1 as described in Algorithm 1 using a learning rate 5e-6, a mini-batch size $64$ in all cases. The training and validation set are same as in Phase 1. Similar to [51], we collect 256 sentences at once and adopt 4 inner epochs when doing the policy optimization. The clipping rate $\epsilon$ is set to 0.2, and the max length of the model is 512. We use temperature = 1 when sampling from $q_{\theta}$. We train the model for 5000 steps and select the final checkpoint using the Rouge-L score on the validation set. Our experiments are based on the NVIDIA V100 32G GPUs. Distilling LLaMA-7B from LLaMA-13B takes less than 10 ours on 16 GPUs.
@@ B.2 Automatic Evaluation Details
During the evaluation, we sample the responses from each model using temperature = 1, a max-length limit of 512, and random seeds [10, 20, 30, 40, 50]. Similar to [67], we adopt a prompt wrapper shown in Figure 9 to convert each instruction-response pair to a sentence. For the GPT-4 feedback, we apply the prompt in Figure 10 and set the temperature = 0.7. For the classification tasks in the “Calibration” paragraph of Section 3.3, we prompt the model to do zero-shot text classification with the templates in Figure 11 and 12.
@@ B.3 Human Evaluation Details
Following [53], we use SelfInst [74] to perform human evaluation. We randomly sampled 50 prompts because we found that more prompts do not affect the results much. We ask the annotators to compare the responses generated by the baseline models with MiniLLM and decide which response is preferred or neither of them is significantly better. Note that which model the responses come from is invisible to the annotators. The interface presented to annotators is shown in Figure 13.
@@ B.4 Details About Generation Diversity Metrics
  • “Dist-4” is a fraction: $N/C$ , where $N$ is the number of the distinct 4-grams in the generated responses and $C$ is the total number of 4-grams. It is a widely used metric to measure the generation diversity of a language model [37]. The $(N/C)$s on the Dolly test set across 5 random seeds are shown in Table 5. Table 7 reports the average values across the 5 random seeds.
  • “Loss” is the negative log-likelihood loss on the test set $\mathcal{D}_{\text{Test}}$: $-\sum_{\bm{x},\bm{y}\sim\mathcal{D}_{\text{Test}}}\log q_{\theta}(\bm{y}|\bm{x})$. It measures the mode coverage of the real data distribution because it is essentially the forward KLD between the real data distribution and the model output distribution. This relates to diversity as in the ability to generate different generations given one context with different random seeds.
@@ B.5 Exposure Bias Analysis
Following [2], we compute the ExAccErr with the following formula:
@@ C.1 GPT-J as the Teacher Model
We present the evaluation results when using GPT-J as the teacher model and GPT-2-760M, GPT-2-1.5B, and GPT-Neo-2.7B [6] as the student models in Table 6. MiniLLM outperforms the baselines in most cases.
@@ C.4 More Ablation Studies
In Table 16, we study the effect of adding the pre-training loss in Algorithm 1 by comparing MiniLLM with its variant where the language modeling loss on the pre-training corpus is removed (w/o $\mathcal{L}_{\text{PT}}$). We have a similar observation as [51] that adding the pre-training loss helps to preserve the abilities on canonical NLP tasks while keeping the performance on instruction-following tasks nearly unchanged.
@@ Appendix D Cases
Instruction
During writing, we added an asterisk for the word that did not come to mind. You will need to provide several examples to demonstrate all the words that can be used in the sentence instead of the asterisk.
Input
we * to know
SFT w/o KD
We can use “know” in the sentence instead of “we * to know”
KD
@@
Instruction
Take the title of the product and extract its attributes. The attributes in this case refer to the characteristics of the products, such as their brand and color. Your answer must a list following the format of “attribute: value”.
Input
SAMSUNG 980 PRO SSD 2TB PCIe NVMe Gen 4 Gaming M.2 Internal Solid State Hard Drive Memory Card, Maximum Speed, Thermal Control, MZ-V8P2T0B
SFT w/o KD
@@
KD
attribute: brand, attribute: color, attribute: memory card, attribute: maximum speed, attribute: thermal control
SeqKD
brand: Samsung, color: black, size: 2TB, interface: PCIe NVME Gen 4, design: M.2 Internal

### v5

@@ A.2 Derivation of Equation 2
We compute the gradient of $\mathcal{L}(\theta)=\operatorname{KL}[q_{\theta}||p]$ with respect to $\theta$ using the Policy Gradient Theorem [policy_gradient]:
@@ B.1 Training Details
Our baselines include SFT w/o KD, KD, and SeqKD. For models with less than 1.3B parameters, we search for the learning rates in [5e-4, 1e-4, 5e-5], the batch sizes in [32, 64], and train these models for 20 epochs. For other models, we search for the learning rate in [5e-5, 1e-5, 5e-6], the batch sizes in [32, 64], and train these models for 10 epochs. For KD, we follow [lightpaff] to mix the distillation loss with the language modeling loss on the ground truth responses by a mixture rate of 0.5. The checkpoints of each baseline are selected by the Rouge-L [rouge] scores on the validation set because, as stated in previous works [super-natural-instructions, instruct-gpt], we also find that Rouge-L is better correlated with human judgments.
@@
As stated in Section 2.3, training of MiniLLM has two phases which is similar to Reinforcement Learning from Human Feedback (RLHF;instruct-gpt).
@@
  • Phase 2: We continuously train the model from Phase 1 as described in Algorithm 1 using a learning rate 5e-6, a mini-batch size $64$ in all cases. The training and validation set are same as in Phase 1. Similar to [instruct-gpt], we collect 256 sentences at once and adopt 4 inner epochs when doing the policy optimization. The clipping rate $\epsilon$ is set to 0.2, and the max length of the model is 512. We use temperature = 1 when sampling from $q_{\theta}$. We train the model for 5000 steps and select the final checkpoint using the Rouge-L score on the validation set. Our experiments are based on the NVIDIA V100 32G GPUs. Distilling LLaMA-7B from LLaMA-13B takes less than 10 ours on 16 GPUs.
@@ B.2 Automatic Evaluation Details
During the evaluation, we sample the responses from each model using temperature = 1, a max-length limit of 512, and random seeds [10, 20, 30, 40, 50]. Similar to [alpaca], we adopt a prompt wrapper shown in Figure 9 to convert each instruction-response pair to a sentence. For the GPT-4 feedback, we apply the prompt in Figure 10 and set the temperature = 0.7. For the classification tasks in the “Calibration” paragraph of Section 3.3, we prompt the model to do zero-shot text classification with the templates in Figure 11 and 12.
@@ B.3 Human Evaluation Details
Following [ITGPT4], we use SelfInst [self_inst] to perform human evaluation. We randomly sampled 50 prompts because we found that more prompts do not affect the results much. We ask the annotators to compare the responses generated by the baseline models with MiniLLM and decide which response is preferred or neither of them is significantly better. Note that which model the responses come from is invisible to the annotators. The interface presented to annotators is shown in Figure 13.
@@ B.4 Details About Generation Diversity Metrics
  • “Dist-4” is a fraction: $N/C$ , where $N$ is the number of the distinct 4-grams in the generated responses and $C$ is the total number of 4-grams. It is a widely used metric to measure the generation diversity of a language model [dist4]. The $(N/C)$s on the Dolly test set across 5 random seeds are shown in Table 5. Table 7 reports the average values across the 5 random seeds.
  • “Loss” is the negative log-likelihood loss on the test set $\mathcal{D}_{\text{Test}}$: $\sum_{\bm{x},\bm{y}\sim\mathcal{D}_{\text{Test}}}\log q_{\theta}(\bm{y}|\bm{x})$. It measures the mode coverage of the real data distribution because it is essentially the forward KLD between the real data distribution and the model output distribution. This relates to diversity as in the ability to generate different generations given one context with different random seeds.
@@ B.5 Exposure Bias Analysis
Following [eb_measure], we compute the ExAccErr with the following formula:
@@ C.1 GPT-J as the Teacher Model
We present the evaluation results when using GPT-J as the teacher model and GPT-2-760M, GPT-2-1.5B, and GPT-Neo-2.7B [gpt-neo] as the student models in Table 6. MiniLLM outperforms the baselines in most cases.
@@ C.4 More Ablation Studies
In Table 16, we study the effect of adding the pre-training loss in Algorithm 1 by comparing MiniLLM with its variant where the language modeling loss on the pre-training corpus is removed (w/o $\mathcal{L}_{\text{PT}}$). We have a similar observation as [instruct-gpt] that adding the pre-training loss helps to preserve the abilities on canonical NLP tasks while keeping the performance on instruction-following tasks nearly unchanged.
@@ Appendix D Cases
Instruction During writing, we added an asterisk for the word that did not come to mind. You will need to provide several examples to demonstrate all the words that can be used in the sentence instead of the asterisk.
Input we * to know
SFT w/o KD We can use “know” in the sentence instead of “we * to know”
@@
Instruction Take the title of the product and extract its attributes. The attributes in this case refer to the characteristics of the products, such as their brand and color. Your answer must a list following the format of “attribute: value”.
Input SAMSUNG 980 PRO SSD 2TB PCIe NVMe Gen 4 Gaming M.2 Internal Solid State Hard Drive Memory Card, Maximum Speed, Thermal Control, MZ-V8P2T0B
@@
KD attribute: brand, attribute: color, attribute: memory card, attribute: maximum speed, attribute: thermal control
SeqKD brand: Samsung, color: black, size: 2TB, interface: PCIe NVME Gen 4, design: M.2 Internal
[…]